OpenStudy (anonymous):

Divide. (–6m9 – 6m8 – 16m6) ÷ (2m3)

OpenStudy (callisto):

-3 m^6 -3m^5-8m^3 -------------------- 2m^3 )–6m^9 – 6m^8 – 16m^6 -6m^9 --------------------- -6m^8-16m^6 -6m^8 ---------------- -16m^6 -16m^6 -----------

OpenStudy (anonymous):

Here is another way. If you wanted to divide a number say 460 by 20, you can first divide it by 10 (getting 46) and then divide it by 2 (getting 23). Do this one in the same way. First divide it by 2. The divide again by m^3.

OpenStudy (anonymous):

–3m9 – 3m8 – 8m6 –3m6 – 6m8 – 16m3 –3m6 – 3m5 – 8m3 –3m6 – 3m5 – 16m3

OpenStudy (callisto):

Ah... that reminds me something :P \[(–6m^9 – 6m^8 – 16m^6) ÷ (2m^3)\]\[ = –6m^9 ÷ (2m^3) – 6m^8 ÷ (2m^3)– 16m^6) ÷ (2m^3)\]\[ = (–6/2)m^{9-3} – (6/2)m^{8 -3}– (16/2)m^{6-3}\]\[ = (–3)m^{6} – (3)m^{5}– (8)m^{3}\]

OpenStudy (anonymous):

thank you again

OpenStudy (callisto):

Welcome :)